Dimension Reduction for Functionals on Solenoidal Vector Fields

Stefan Krömer Mathematisches Institut Universitaet zu Koeln D-50923 Koeln Germany skroemer@math.uni-koeln.de

Abstract: We study integral functionals constrained to divergence-free vector fields in L^{p} on a thin domain, under standard p-growth and coercivity assumptions, 1 < p < ∞. We prove that as the thickness of the domain goes to zero, the Gamma-limit with respect to weak convergence in L^{p} is always given by the associated functional with convexified energy density wherever it is finite. Remarkably, this happens despite the fact that relaxation of nonconvex functionals subject to the limiting constraint can give rise to a nonlocal functional as illustrated in an example.